### v1

@@ III-E Truthfulness
Truthfulness pertains to the factual accuracy of information or statements, specifically their alignment with real-world facts [230]. Large language models (LLMs), trained on extensive textual datasets, may occasionally blur the distinction between fact and fiction in their responses [122, 123], potentially undermining user trust. A prominent issue related to truthfulness is hallucination [81, 119, 120], wherein models generate responses that seem plausible but are factually incorrect, thus misleading users. For example, when prompted about the “Four Great Inventions” of ancient China, a model might inaccurately list “gunpowder, the compass, silk, and porcelain,” while the correct answer is “papermaking, the compass, gunpowder, and printing.”
@@
AI-LIEDAR: [26] aims to investigate the balance between utility [118] and truthfulness [117] in LLMs. These models may produce inaccurate information to meet user demands, particularly when task objectives conflict with truthful disclosure, as in promoting defective products. While previous studies have mainly addressed "hallucinations" (i.e., the generation of unsupported content), there has been limited focus on evaluating model truthfulness in scenarios where user instructions are ambiguous and prioritize utility. The AI-LIEDAR framework includes 60 real-life-inspired multi-turn dialogue scenarios, each shaped by one of three motivations for deception: benefits, public image, and emotion. Using the Sotopia platform [121], the framework simulates interactions between users and models, with a psychology-based truth detector classifying responses as fully truthful, partially truthful (e.g., through omission), or entirely false. Additionally, the study investigates model steerability, testing whether prompts can effectively direct responses toward truthfulness or deception. Findings reveal that LLMs rarely sustain complete truthfulness throughout interactions, with an overall truthfulness rate below 50%. Notably, a negative correlation emerges between truthfulness and goal achievement, especially in scenarios involving objective metrics, such as product promotion. Even when models are explicitly prompted toward “truthful” or “deceptive” responses, their truthfulness levels remain variable. The authors advocate for the development of more refined evaluation frameworks and robust guidance mechanisms to enhance model reliability across diverse contexts.
THaMES: [127] is a framework dedicated to evaluating and mitigating hallucinations generated by LLMs. This study is motivated by the frequent emergence of inaccurate information, especially in complex, domain-specific texts generated by existing models. The THaMES framework comprises three primary modules—QA set generation, hallucination benchmarking, and mitigation strategies—designed to identify and reduce hallucinatory outputs effectively. The framework begins by constructing high-quality question-answer test sets through weighted sampling and advanced question-generation techniques, enabling a thorough evaluation of hallucination instances. Next, multiple benchmark metrics are employed to assess models’ capabilities in both recognizing and generating hallucinatory content. Following this, THaMES implements a series of refined mitigation strategies, including In-Context Learning (ICL) [124], Retrieval-Augmented Generation (RAG) [125], and parameter-efficient fine-tuning (e.g., PEFT) [126, 128], aimed at enhancing model performance and reducing hallucinations. To support diverse model architectures, THaMES offers a variety of mitigation options, allowing users to select the most suitable approach based on specific model characteristics. Experimental results reveal that the commercial model GPT-4 showed significant improvement under the RAG strategy, which leverages external knowledge integration, while open-weight models, such as Llama-3.1, achieved higher reasoning accuracy through ICL. Additionally, PEFT fine-tuning applied to Llama-3.1 demonstrated a marked reduction in hallucinations. By elucidating the effects of different mitigation strategies on hallucination reduction across various models, the THaMES framework establishes a new standard and provides valuable guidance for the ongoing development of LLMs.
FACTKB: [129] addresses the challenge of evaluating truthful accuracy in automated summarization [130, 131]. Current generative models often produce inaccurate factual information, particularly with respect to entities and relationships [135]. To enable more generalized and robust truthfulness evaluation, FACTKB combines pretrained language models with knowledge bases, employing three entity-centered pretraining strategies—Entity Wiki, Evidence Extraction, and Knowledge Walk—to enhance the model’s representation of entities and relationships. In detail, Entity Wiki enriches the model’s understanding of specific entities by integrating relevant background knowledge directly from the knowledge base. Evidence Extraction extracts supporting evidence from auxiliary information linked to entities, enhancing the model’s capacity to verify facts within context. Finally, Knowledge Walk leverages multi-hop knowledge paths within the knowledge base, strengthening the model’s compositional reasoning ability for complex relationships. By training with these strategies, the language model improves its evaluation of entity and relationship accuracy, enabling fine-tuning for precise factual error detection. Experimental results show that FACTKB surpasses other methods in identifying semantic framework errors, particularly those related to entities and relationships, while minimizing preprocessing requirements, thereby making it adaptable to diverse datasets. FACTKB not only aligns well with human judgment but also achieves high accuracy in detecting various error types, especially in complex domains. Ultimately, this framework offers an accurate and accessible method for truthfulness evaluation, setting a reliable standard for fact-checking in text generation across multiple domains.
ReEval: [146] seeks to create new evaluation datasets using adversarial attacks [144, 145] to test whether retrieval-augmented LLMs [142, 143] accurately reference provided evidence in their responses. Traditional static question-answering datasets, frequently included in LLM pretraining data, often lead models to rely on memorization rather than genuine evidence, complicating objective assessments of model reliability. ReEval is motivated by the need for dynamic data generation that can automatically detect hallucinations—instances where model outputs deviate from the actual evidence. ReEval incorporates two data generation techniques: answer swapping and context enriching. In answer swapping, portions of the supporting evidence containing the answer are altered to examine whether the model aligns with the new answer context. In context enriching, supplementary information is added to the original context, enhancing question complexity while preserving the core evidence. Experiments were conducted on multiple open-domain question-answering datasets using LLMs, including GPT-4 and ChatGPT. Results indicate that all models exhibited a significant accuracy decline on adversarial test data, despite varied prompts, suggesting that ReEval effectively induced hallucination phenomena. Furthermore, ReEval’s adversarial test samples demonstrated transferability; samples generated by smaller models were applicable to larger models, substantially reducing evaluation costs. Additionally, human reviewers rated the ReEval-generated test data positively in terms of readability and evidential support, confirming that these datasets realistically capture complex real-world scenarios.
HaluEval: [147] addresses the prevalent issue of hallucinations in LLMs, where models often generate unverifiable or inaccurate information. To systematically evaluate LLMs’ capabilities in hallucination detection, HaluEval introduces a dataset of 35,000 samples covering three tasks: knowledge-based dialogue [149], question answering [148], and text summarization [150], with each sample either manually annotated or automatically generated. By employing a two-step framework involving sampling and filtering, HaluEval ensures both diversity and high complexity in its samples, aiming to rigorously test models’ hallucination recognition capabilities. The dataset is constructed by first sampling user queries from various public sources, with ChatGPT generating responses likely to contain hallucinations. Responses with low similarity, as calculated using BERTScore [151], are then filtered and annotated for hallucinated content. Additionally, HaluEval incorporates task-specific, automatically generated samples through diverse sampling and filtering methods to create a challenging set of hallucination cases. Experimental results reveal notable shortcomings in current LLMs’ hallucination detection abilities. For example, ChatGPT achieved an accuracy of only 58.53% in identifying hallucinations within text summaries. The study further indicates that incorporating external knowledge or adding reasoning steps significantly enhances hallucination detection for certain tasks, though these improvements are less pronounced in dialogue contexts. Additionally, HaluEval reveals that hallucination frequency is strongly topic-dependent, with higher rates observed in content related to film, technology, and climate. This work underscores critical blind spots in LLMs’ performance on specific topics and complex dialogues, offering essential insights for developing more reliable models in the future.
In addition to the studies mentioned above, other researchers have also explored methods for evaluating the truthfulness of LLMs. Wang et al. [152] proposed OpenFactCheck, a modular framework that provides a unified system for evaluating the factual accuracy of content generated by large language models (LLMs). OpenFactCheck integrates multiple evaluation modules, allowing customizable truthfulness assessments. Demonstrating high consistency across diverse datasets, particularly in dynamic information domains, its adversarially generated samples effectively test factual verification and induce hallucination phenomena in complex knowledge areas, making it ideal for assessing LLM reliability. GraphEval [153] detects hallucinations in LLM outputs by constructing knowledge graphs (KGs) [154] in two steps: KG construction and triple consistency verification. By limiting LLM calls to a single KG construction pass, this approach reduces computational costs and enhances interpretability by pinpointing inconsistencies. Combined with natural language inference (NLI) models, GraphEval significantly improves balanced accuracy in hallucination detection, boosting NLI performance by approximately 6.2 points on benchmarks like SummEval [155] and QAGS-C [156], especially for longer outputs. Zhu et al. [157] introduced KG-FPQ, a benchmark that leverages knowledge graphs to generate False Premise Questions (FPQs) for assessing truthfulness hallucinations in LLMs. Covering art, people, and places, the KG-FPQ dataset includes around 178,000 FPQs, enabling comprehensive tests of hallucination handling in both discriminative and generative tasks [158, 159, 160]. The framework also introduces FPQ-Judge, an automated hallucination evaluator. Findings reveal LLM susceptibility to hallucinations under false premises, with FPQ complexity and task format significantly impacting hallucination performance, underscoring the need for improved truthfulness across varied tasks and domains.
@@ III-F Other
In addition to commonly studied evaluation dimensions—such as toxicity, robustness, ethics, bias and fairness, and truthfulness—LLMs require rigorous evaluation across several other critical areas. These include privacy, mental health and medical applications, and code generation, copyright, agents. Assessments in these domains are essential to ensuring that LLMs adhere to standards of safety, legality, and effectiveness, thus supporting their responsible and reliable deployment in real-world contexts.
@@
Privacy: LLMs rely on extensive data for training, posing inherent risks of exposing sensitive user information [161, 162, 163], including personal identifiers, geographic locations, and employment records. Such potential privacy leakage not only erodes user trust but also raises significant data compliance concerns. Accordingly, a systematic assessment of privacy risks in LLMs is critical for identifying and mitigating channels of potential privacy breaches, thereby ensuring model safety and regulatory compliance in practical applications. Li et al. [39] introduced LLM-PBE, a comprehensive toolkit for evaluating privacy risks throughout the LLMs lifecycle. LLM-PBE facilitates privacy assessments across diverse data types, attack vectors, and defense strategies, with a particular focus on understanding how model size, data characteristics, and temporal aspects impact privacy vulnerabilities. The toolkit includes modules for various attack types, such as data extraction [163, 164], membership inference [165], prompt leakage [166], and jailbreaking [167], complemented by defense strategies like differential privacy [168]. Extensive experimentation on datasets such as Enron [169] and ECHR [170] offers a systematic analysis of privacy concerns in LLMs. Additionally, the LLM-PBE framework integrates support for models hosted on platforms such as OpenAI, TogetherAI, and Hugging Face, enabling users to assess privacy risks across various LLMs through API interfaces. Experimental results indicate that larger models are more susceptible to privacy leakage. Additionally, certain data types and placements, such as sensitive information located at the end of documents, are particularly prone to extraction. Furthermore, while differential privacy effectively mitigates leakage risks, it introduces some trade-offs in model performance, underscoring the need to balance privacy protection with practical utility.
Mental Health and Medical Applications: The growing popularity of mental health chatbots, valued for their human-like, context-aware support, has attracted a large user base. However, these chatbots are prone to producing hallucinations or misleading advice and may fail to offer appropriate guidance in crisis situations, such as managing self-harm [171]. Consequently, there is an urgent need for a standardized evaluation framework to assess the safety and reliability of these systems, thereby fostering user trust and facilitating broader adoption in mental health care. Park et al. [34] proposed a dedicated evaluation framework to ensure the safety and reliability of mental health chatbots based on LLMs. Their framework consists of 100 benchmark questions paired with ideal responses, addressing common mental health scenarios. In addition, five guiding questions assist mental health professionals in evaluating models across key dimensions: safety, crisis management, consistency, resource provision, and user autonomy. This framework provides a comprehensive standard for evaluating LLMs in mental health applications, supporting trust-building and the safer integration of these tools into healthcare settings. Similarly, the rapid development of LLMs in medical applications has introduced significant safety concerns. These models, when processing medical information, may generate misleading or unethical recommendations, which pose potential risks to patients and public health. Han et al. [183] introduced MedSafetyBench, a benchmark dataset specifically designed to evaluate the safety of LLMs in medical contexts. MedSafetyBench includes 1,800 examples of harmful medical requests with corresponding safe responses, structured around nine core principles of medical ethics. By generating harmful medical prompts and appropriate responses, the researchers assessed existing LLMs (such as GPT-4 and Llama) to determine their capacity to reject unsafe requests. Experimental results indicate that current medical LLMs frequently fall short of safety standards, often providing unsafe responses to inappropriate prompts. The researchers also conducted fine-tuning experiments using MedSafetyBench, which demonstrated that fine-tuning with this dataset significantly enhances model safety while maintaining the accuracy of medical knowledge.
Code Generation: With the growing application of LLMs in programming assistance, their potential to generate insecure code or respond to malicious prompts introduces substantial cybersecurity risks [172, 173, 174]. For example, LLMs may produce code vulnerable to common attacks, such as SQL injection or cross-site scripting, leaving generated applications susceptible to exploitation. Consequently, assessing the security of LLMs in code generation has become imperative. Bhatt et al. [36] introduced the CYBERSECEVAL framework, designed to evaluate the security of LLM-generated code and their responsiveness to cyberattack-related prompts. CYBERSECEVAL includes two main evaluation components: insecure coding tests and cyberattack helpfulness tests. The insecure coding tests utilize a “vulnerability detector” to generate cases reflecting common programming weaknesses, assessing whether models generate insecure code through both code completion and prompt-based tasks. The cyberattack helpfulness tests involve researcher-designed, attack-related prompts to evaluate if model responses might facilitate malicious actions. Experimental findings indicate that all tested models produced insecure code in 30% of insecure coding scenarios, with models possessing stronger programming capabilities posing greater security risks. Additionally, an average of 53% of cyberattack prompts elicited responses aligning with attack objectives, underscoring LLMs’ limitations in handling malicious requests. The study also noted that models with more advanced programming capabilities demonstrated a higher level of compliance in cyberattack scenarios. CYBERSECEVAL thus provides a systematic tool for evaluating LLMs cybersecurity, supporting the development of safer AI systems. Additionally, Wang et al. [35] developed the CodeSecEval dataset to evaluate the security of LLMs in code generation and repair. CodeSecEval comprises 180 samples across 44 critical vulnerability types and is organized into two subsets (SecEvalBase and SecEvalPlus), designed for assessing secure code generation and repair. Their study conducted multiple experiments to examine LLMs’ secure code generation abilities and proposed several enhancement strategies, including “vulnerability-aware information” (explicitly indicating vulnerability risks) and “insecure code explanations” (providing explanations for vulnerabilities) to improve LLMs security.
@@
Copyright: A systematic approach to copyright assessment is essential for effectively identifying and preventing the generation of infringing content, thus ensuring compliance and facilitating the safe, lawful use of LLMs across various creative domains. This approach mitigates potential legal risks related to copyright for both companies and users. Liu et al. [37] introduced the SHIELD framework, a tool designed to evaluate and prevent copyright violations by LLMs during text generation [175]. As LLMs find broader application in content creation, concerns about copyright infringement have become more pronounced, particularly due to their potential to generate protected content without authorization [176]. This study addresses gaps in existing copyright protection mechanisms, which may be overly restrictive or insufficiently responsive to unauthorized content generation. The SHIELD framework employs a curated copyright dataset to evaluate whether model-generated text adheres to copyright standards and introduces a lightweight, real-time proxy defense mechanism to dynamically detect and block potentially infringing content. The SHIELD framework consists of three primary modules: the Copyright Material Detector, Copyright Status Verifier, and Copyright Status Guide. The Copyright Material Detector uses an N-gram language model to identify copyrighted segments within generated content; the Copyright Status Verifier performs real-time online checks to verify the copyright status of content; and the Copyright Status Guide provides contextual guidance to direct the model toward producing copyright-safe content. This defense mechanism is designed to quickly identify and reject sensitive requests while maintaining flexibility in generating public domain text. Experimental results demonstrate that the SHIELD framework significantly reduces the generation of copyrighted material. It also achieves higher refusal rates in response to “jailbreak attacks” [45, 104, 177] (attempts to bypass model safeguards), underscoring its robustness. Furthermore, SHIELD effectively avoids over-protection by not interfering with the generation of public domain content, offering a practical and reliable solution for copyright compliance.
Agents: LLMs Agents serve as automated systems powered by LLMs, designed to analyze user input continuously and provide outputs that meet specific objectives, facilitating complex operations and interactive task completion[178, 179]. Due to their application in diverse, intricate scenarios, rigorous evaluation is crucial to optimize performance and ensure safety. Yuan et al.[40] introduced R-Judge, a benchmark dataset designed to assess LLMs’ awareness of safety risks in multi-turn interactions. R-Judge comprises 569 multi-turn records across 27 risk scenarios in programming, IoT, software, networking, and finance, focusing on ten risk types like privacy breaches and property loss. This benchmark reveals current models’ limitations in multi-dimensional safety risk handling, emphasizing the need for diverse, high-quality data to enhance LLM safety. Similarly, Zhu et al.[41] proposed RiskAwareBench, a framework to evaluate physical risk awareness in LLM-based agents, where “physical risk” refers to potential harm to property or individuals from decisions by agents such as robots. This framework utilizes the PhysicalRisk dataset, which includes safety guidelines and scenarios to systematically test LLMs’ capacity to recognize and mitigate physical risks. Zhang et al.[42] developed Agent Security Bench (ASB) to assess the security resilience of LLMs agents against security threats, including prompt injection, memory poisoning, and backdoor attacks [180, 181, 182]. Covering ten domains, such as e-commerce and autonomous driving, ASB evaluates ten agents using over 400 tools and tasks across 23 attack and defense strategies, revealing an 84.3% average success rate for mixed attacks. These findings highlight significant vulnerabilities in current LLMs agents and the limited effectiveness of existing defenses against complex attacks.

### v2

@@ III-E Truthfulness
Truthfulness refers to the factual accuracy of information generated by LLMs, specifically their alignment with real-world facts [146]. Ensuring truthfulness is crucial for building user trust and preventing the dissemination of misleading information. A central challenge in this domain is the phenomenon of hallucination [147, 148, 149], where LLMs produce seemingly plausible but factually inaccurate or contextually disconnected responses. Consequently, most existing research on truthfulness evaluation has focused on hallucination [150, 151, 152], while a smaller body of work examines other dimensions such as utility-driven deception.
III-E1 Hallucination Evaluation
A substantial body of research has developed evaluation methods and benchmarks to systematically assess hallucinations in LLMs. One stream emphasizes dataset construction under realistic interaction scenarios. For example, ReEval [50] create new evaluation datasets using adversarial attacks to test whether retrieval-augmented LLMs accurately reference provided evidence in their responses. ReEval incorporates two data generation techniques: answer swapping and context enriching. In answer swapping, portions of the supporting evidence containing the answer are altered to examine whether the model aligns with the new answer context. In context enriching, supplementary information is added to the original context, enhancing question complexity while preserving the core evidence. Results show that all models exhibited a significant hallucination on adversarial test data. Moreover, HaluEval [51] provides a large-scale benchmark that evaluate LLMs’ capabilities in hallucination detection, which cover knowledge-based dialogue, question answering, and text summarization. Experimental results reveal notable shortcomings in current LLMs’ hallucination detection abilities, and hallucination frequency is strongly topic-dependent, with higher rates observed in content related to film, technology, and climate.
Another stream leverages structured external knowledge to support factual verification. GraphEval [53] detects hallucinations in LLMs’ outputs by constructing knowledge graphs (KGs) in two steps: KG construction and triple consistency verification. By limiting LLM calls to a single KG construction pass, this approach reduces computational costs and enhances interpretability by pinpointing inconsistencies. Combined with natural language inference (NLI) models, GraphEval significantly improves balanced accuracy in hallucination detection, boosting NLI performance by approximately 6.2 points on benchmarks like SummEval [153] and QAGS-C [154], especially for longer outputs. Similarly, KG-FPQ [54] leverages KGs to generate False Premise Questions (FPQs) for assessing truthfulness hallucinations in LLMs. By selecting entities from art, people, and places, KG-FPQ first extracts real triplets from the KGs and generates false premises through edit true triplets to false triplets. Then, it utilizes ChatGPT’s natural language generation capability to construct semantically rich FPQs. Experiments show that when edited object has a closer distance with the subject or has a stronger association with the original object, FPQs are more confusing to LLMs. And LLMs perform worse at generating factual statements than at distinguishing them when faced with FPQs.
In addition, some works propose unified framework for the evaluation of hallucinations. THaMES [49] introduces a end-to-end tool for hallucination detection and mitigation, which comprises three primary modules—QA set generation, hallucination benchmarking, and mitigation strategies. The framework begins by constructing high-quality question-answer test sets through weighted sampling and advanced question-generation techniques. Next, multiple metrics from RAGAS [155] and HaluEval [51] are used to evaluate hallucination. Following this, THaMES implements a series of refined mitigation strategies, including In-Context Learning (ICL) [156], Retrieval-Augmented Generation (RAG) [157], and parameter-efficient fine-tuning (e.g., PEFT) [158, 159]. In order to solve the problem of difficulty in free-form response evaluation in open-domain and fragmentation of datasets and metrics, Iqbal et al. [52] propose OpenFactCheck. The framework consists of three collaborative modules: First, RESPONSEEVAL supports users to customize the fact-checker, integrating the claim processors, evidence retrievers and verifiers of existing systems [160, 161, 162] to realize factual error detection in a free-form document; Second, LLMEVAL builds a FactQA benchmark, which divides factual deviations into three categories: knowledge errors, over-commitment errors, and disability errors, and unifies evaluation metrics (such as exact matching, FreshEval [163], and the proportion of true statements) to systematically evaluate LLM’s factual capabilities in multiple dimensions; Third, CHECKEREVAL assesses the verification accuracy of fact-checkers, equipped with a leaderboard in terms of accuracy, latency, and costs.
III-E2 Beyond Hallucinations
@@
While hallucination has attracted primary attention, other dimensions of truthfulness are also emerging. AI-LIEDAR [48] aims to investigate the balance between utility [164] and truthfulness [165] in LLMs. These models may produce inaccurate information to meet user demands, particularly when task objectives conflict with truthful disclosure, as in promoting defective products. The AI-LIEDAR framework includes 60 real-life-inspired multi-turn dialogue scenarios, each shaped by one of three motivations for deception: benefits, public image, and emotion. Using the Sotopia platform [166], the framework simulates interactions between users and models, with a psychology-based truth detector classifying responses as fully truthful, partially truthful (e.g., through omission), or entirely false. Additionally, the study investigates model steerability, testing whether prompts can effectively direct responses toward truthfulness or deception. Findings reveal that LLMs rarely sustain complete truthfulness throughout interactions, with an overall truthfulness rate below 50%. Notably, a negative correlation emerges between truthfulness and goal achievement, especially in scenarios involving objective metrics, such as product promotion. Even when models are explicitly prompted toward “truthful” or “deceptive” responses, their truthfulness levels remain variable.
Despite notable advances, truthfulness evaluation remains constrained by several challenges. Current evaluation metrics for hallucinations face significant limitations, including a lack of universal applicability, inadequate comprehensiveness, and vulnerability to dataset leakage [167], which undermine their robustness and fairness. Furthermore, many recent efforts have turned to the paradigm of LLM-as-a-Judge [145] to assess truthfulness. While this approach offers scalability and flexibility, it also introduces a recursive dilemma: if evaluating one model requires relying on a stronger and more reliable model, the trustworthiness of the evaluation itself becomes questionable. These challenges highlight the urgent need for innovative and standardized evaluation systems that can holistically capture different forms of untruthfulness, while reducing dependency on opaque or unverifiable evaluators.
@@ III-F Other
In addition to commonly studied evaluation dimensions—such as toxicity, robustness, ethics, bias and fairness, and truthfulness—LLMs require rigorous evaluation across several other critical areas. These include privacy, mental health and medical applications, and code generation, copyright. Assessments in these domains are essential to ensuring that LLMs adhere to standards of safety, legality, and effectiveness, thus supporting their responsible and reliable deployment in real-world contexts.
III-F1 Privacy
@@
LLMs rely on extensive data for training, posing inherent risks of leaking sensitive user information [168, 169, 170], including personal identifiers, geographic locations, and employment-related information. Consequently, a systematic assessment of privacy risks in LLMs is critical to identify and mitigate potential breach channels, thereby ensuring model safety and compliance in real-world applications. Li et al. [55] introduced LLM-PBE, a comprehensive toolkit for evaluating privacy risks throughout the LLM lifecycle. LLM-PBE provides systematic assessments across diverse data types, multiple attack vectors, and corresponding defense strategies, with particular emphasis on analyzing how model size, data characteristics, and temporal factors impact privacy vulnerabilities. The toolkit incorporates modules for different attack types, such as data extraction [170, 171], membership inference [172], prompt leakage [173], and jailbreaking [95], complemented by defense methods like differential privacy [174]. Moreover, the framework integrates support for models hosted on platforms such as OpenAI, TogetherAI, and Hugging Face, enabling privacy assessments through API interfaces. Extensive experimentation on datasets such as Enron [175] and ECHR [176] enables a systematic analysis of privacy vulnerabilities in LLMs. Empirical results demonstrate that larger models exhibit higher susceptibility to privacy leakage. Certain data types and placements, such as sensitive information located at the end of documents, are more vulnerable to extraction. Furthermore, while differential privacy effectively mitigates leakage risks, it introduces notable trade-offs in model utility, highlighting the necessity of balancing privacy preservation with performance. Despite these advances, key challenges remain. Privacy attacks evolve rapidly, demanding adaptive evaluation frameworks that can capture emerging threats. The trade-off between privacy preservation and model effectiveness remains unresolved, particularly in high-stakes domains. Addressing these challenges is essential for building trustworthy and privacy-preserving LLMs.
III-F2 Copyright
@@
A systematic approach to copyright assessment is essential for systematically identifying and mitigating the generation of infringing content, thereby ensuring compliance and enabling the safe and lawful use of LLMs across diverse creative domains. As LLMs are increasingly applied in content creation, concerns regarding copyright infringement have intensified, particularly due to their ability to reproduce protected content without authorization [177]. Existing copyright protection mechanisms often prove either overly restrictive or insufficiently adaptive, leaving gaps in safeguarding against unauthorized content generation. Liu et al. [60] introduced the SHIELD framework, a tool designed to evaluate and prevent copyright violations by LLMs during text generation [178]. SHIELD employs a curated copyright dataset to assess whether generated text adheres to copyright standards and incorporates a lightweight and real-time proxy-based defense mechanism to dynamically detect and block potentially infringing content. The framework consists of three primary modules: the Copyright Material Detector, which uses an N-gram model to identify copyrighted segments; the Copyright Status Verifier, which performs real-time checks to verify copyright status; and the Copyright Status Guide, which provides contextual feedback to direct the model toward producing copyright-safe outputs. This mechanism is designed to promptly detect and reject infringing requests while maintaining flexibility in generating public domain content. However, balancing copyright protection with creative flexibility is a persistent issue, as overly strict safeguards may compromise usability. Jurisdictional diversity and the evolving nature of copyright status further complicate the establishment of standardized evaluation practices. Moreover, the scarcity of large-scale, high-quality copyright datasets limits the creation of comprehensive benchmarks. Addressing these challenges will be critical for ensuring that LLMs can be responsibly applied in creative and commercial domains.
III-F3 Code Generation
Unlike harmful content in other domains that directly affects individuals or society, risks in code generation stem from insecure code or code that facilitates cyberattacks, posing significant cybersecurity threats [179, 180, 181]. For instance, LLMs may produce code vulnerable to common exploits such as SQL injection or cross-site scripting, rendering applications susceptible to compromise. Evaluating the security of LLMs in code generation has therefore become a critical research priority. Bhatt et al. [57] introduced the CYBERSECEVAL framework, which evaluates the security of LLM-generated code and their susceptibility to cyberattack-related prompts. It includes two components: insecure coding tests, which employ a vulnerability detector to construct cases reflecting common programming weaknesses, and cyberattack helpfulness tests, which use attack-related prompts to assess whether model responses facilitate malicious actions. Empirical findings reveal that all tested models generated insecure code in about 30% of insecure coding scenarios, and that models with stronger programming capabilities posed greater security risks. Furthermore, approximately 53% of cyberattack prompts elicited responses aligned with adversarial objectives, highlighting LLMs’ limitations in handling malicious requests. CYBERSECEVAL thus offers a systematic framework for evaluating LLM cybersecurity. In addition, Peng et al. [58] identified limitations in existing benchmarks such as CYBERSECEVAL, including vague specifications, the separation of functional and security evaluations, and instability in static analysis due to high false positive and false negative rates. To address these issues, they proposed the CWEVAL framework and the CWEVAL-BENCH benchmark. CWEVAL introduces a result-driven synchronous evaluation mechanism that replaces static with dynamic analysis, monitoring runtime attributes such as execution time, memory access, and data integrity to verify both functional correctness and security simultaneously. CWEVAL-BENCH comprises 119 security-critical tasks spanning 31 CWE vulnerability types and five programming languages, with secure and insecure reference implementations to ensure reproducibility. Experimental results indicate that about 30% of functionally correct code generated by mainstream LLMs still contains vulnerabilities. Larger models demonstrated comparatively stronger security capabilities, while simple security prompts improved performance, though fine-tuning sometimes degraded functional correctness. Li et al. [59] further proposed the SafeGenBench framework and its automated evaluation system to overcome limited scenario coverage and single-method evaluation in existing benchmarks. SafeGenBench incorporates 558 test cases covering 13 programming languages and 44 CWE vulnerabilities, integrating authoritative standards such as OWASP Top-10 and CWE Top-25. Its design simulates realistic development scenarios by deliberately avoiding explicit security terms, ensuring alignment with practical usage. Its key innovation lies in a dual-layer evaluation mechanism: Static Application Security Testing (SAST) for broad-spectrum vulnerability scanning, complemented by an LLM-based judge for deep semantic analysis of task-specific vulnerabilities. These complementary methods mitigate the shortcomings of relying on a single approach. Looking ahead, the evaluation of code generation security faces persistent challenges. Software vulnerabilities continue to evolve rapidly, requiring benchmarks and frameworks that can adapt to emerging threats. At the same time, achieving accurate, efficient, and user-friendly evaluation remains a difficult task. Addressing these issues will be essential to ensure that LLM-assisted software development remains both secure and reliable in real-world settings.
III-F4 Downstream Applications in Medicine, Finance, and Law
As LLMs are increasingly adopted in high-stakes fields such as medicine, finance, and law, systematic evaluation of their domain-specific safety has become increasingly critical. Han et al. [56] introduced MedSafetyBench, a benchmark dataset specifically designed to assess LLM safety in medical contexts. MedSafetyBench comprises 1,800 harmful medical requests paired with safe responses, organized around nine fundamental principles of medical ethics. By generating harmful prompts and corresponding responses, the study evaluated models such as GPT-4 and Llama to assess their ability to reject unsafe requests. Results indicate that current medical LLMs often fail to meet established safety standards, producing unsafe responses when confronted with inappropriate prompts. Hui et al. [61] introduced Trident-Bench, the first benchmark targeting LLM safety across finance, medicine, and law. Grounded in the AMA Principles of Medical Ethics, the ABA Model Rules of Professional Conduct, and the CFA Institute Code of Ethics, Trident-Bench contains 2,652 harmful prompts with expert-validated safe responses. This benchmark addresses gaps left by prior evaluations that emphasized accuracy and domain knowledge while overlooking professional ethical compliance. An empirical study of 19 general-purpose, domain-specific, and safety-aligned models revealed critical safety gaps: while advanced closed-source models such as GPT-4o were able to meet basic safety expectations, domain-specific models frequently struggled with subtle ethical issues. Safety-aligned models, such as Llama Guard, performed better, indicating that domain specialization does not necessarily ensure ethical robustness. Although these efforts mark important progress, research on safety evaluation in medicine, finance, and law remains limited and lags behind the rapid deployment of LLMs in these domains. The specialized knowledge required in each field makes the design of rigorous and representative benchmarks challenging. Furthermore, variations in ethical norms and legal regulations across jurisdictions complicate the establishment of unified evaluation standards. Addressing these issues is essential for ensuring the responsible and safe integration of LLMs into high-stakes professional practices. At the same time, other downstream applications—such as education [182] and transportation [183]—pose comparable safety concerns, highlighting the need to broaden evaluation efforts beyond currently emphasized domains.